OpenStudy (anonymous):

what is the vertex of y^2=4ax-1

OpenStudy (anonymous):

hello amstre. you missed the small latex tutorial

OpenStudy (amistre64):

i seen parts of it i think :)

OpenStudy (amistre64):

power was out for a few hours over here ....

OpenStudy (anonymous):

what is the vertex of y^2=4ax-1

OpenStudy (amistre64):

is that 4a(x-1) ? or just a straight run thru?

OpenStudy (anonymous):

it is 4ax-1

OpenStudy (amistre64):

then you factor out the 4a and get: \[y^2=4a(x-\frac{1}{4a})\]

OpenStudy (amistre64):

and the vertex is at: (1/4a , 0)

OpenStudy (amistre64):

otherwise you have a typo; and you really meant to say: 4a(x-1) in which case its (1,0)

OpenStudy (anonymous):

vertex without a.I want only the numerical value

OpenStudy (amistre64):

you cant have it both ways; either you typed it correctly and we leave it in; or you mistyped and we take it out

OpenStudy (anonymous):

ok Thanks a lot

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

it is \[(\frac{1}{4a},0)\] yes?

OpenStudy (anonymous):

yes absolutely.

OpenStudy (amistre64):

it would be if we get an accurate portrayal of the equation ;)

OpenStudy (anonymous):

whew